OpenStudy (anonymous):

A 10.0 kg ball falling at 10.0 m/s hits a pile of feathers and comes to a complete stop in 10 s. What force does the pile of feathers exert to stop the ball? 10 N 20 N 100 N 200 N I think the answer is 100 N

OpenStudy (ell):

Force(N)=mass(kg) x acceleration(m/s^2) F=ma F=10.0(10.0) F=100N Hope this helped!:)

OpenStudy (anonymous):

awesome thanks :D

OpenStudy (ell):

np!(: